The city of El Centro is approximately 11 square miles in area. If one inch of rain falls on El Centro, how many cubic feet of rain would fall on El Centro? (Hint: convert all units to feet first)
Answer by FrankM(1040)

11 sq miles = 11*5280*5280 or 11*27878400 sq ft. 306,662,400 sq ft.
1 in = 1/12 ft so 306,662,400 * 1/12 = 25,555,200 cubic ft of rain.